Calculate H+ in 0.20 M solution of dichloroacetic acid that contains 0.1 M sodium dichloroacetate. Neglected hydrolysis of sodium salt

Grade:12th pass

1 Answers

Arun
6 years ago
CHCl2COOH CHCl2COO–
 + H+ CHCl2COONa  CHCl2COO–
 + Na+  Ka
 =
2
2
[CHCl COO ][H ]
[CHCl COOH]  
 5 × 10–2 =
(0.1 0.2 )(0.2 )
0.2(1 )
  

 2
1
 =  
  
1
(1 2 )
 1 –  = 2 + 4
2
 4
2
 + 3 –1 = 0   =
8  3 9 16
 = 8  3 5
 = 8
2
 = 0.25
 [H+
] = 0.2  = 0.2 × 0.25 = 0.5